OpenStudy (anonymous):

write the expression in standard form. 7/4-14i

OpenStudy (jdoe0001):

that looks pretty standard to me :|

OpenStudy (anonymous):

lol its not one of the answer choices so its probably not

OpenStudy (jdoe0001):

what choices are they giving? because as far as I undestand it, THAT IS the standard form, the other form 'imaginary' numbers have is trigonometric form, or polar

OpenStudy (anonymous):

-7/53+49/106i thats one

OpenStudy (anonymous):

\[\frac{ -7 }{ 53 }+\frac{ 49 }{ 106 }i\]

OpenStudy (jdoe0001):

ahemm, I got that, BUT, the form is still in standard format, namely \(\color{red}{a+bi}\)

OpenStudy (anonymous):

well it must not be if the question is asking me to put it in standard form

OpenStudy (jdoe0001):

well, I think the wording of the Q is not-so-good, unless they mean they don't want any rationals

OpenStudy (jdoe0001):

you can always get rid of the rational by multiplying by 4

OpenStudy (nubeer):

\[\frac{ 7 }{ 4-14i }\] is this the question?

OpenStudy (anonymous):

Yes

OpenStudy (nubeer):

well its not standard form.. u have to convert it and u can do that by doing.

OpenStudy (nubeer):

\[\frac{ 7 }{ 4-14i }\times \frac{ 4+14i }{ 4+14i }\] now solve this.

OpenStudy (anonymous):

Well the denominator just cancels out right?

OpenStudy (nubeer):

no u dont have to cancel here u have to multiply here.. its called rationalizing.

OpenStudy (anonymous):

My class is online i rely on this site to teach me, I have no teacher

OpenStudy (nubeer):

its ok.. just telling u what i amd doing.. :)

OpenStudy (anonymous):

lol thanks

OpenStudy (nubeer):

so what i am doing is called rationalizing.. can u multiply what i wrote.. don't cancel the 4+14i

OpenStudy (anonymous):

multiply straight across

OpenStudy (nubeer):

yes... n tell me what u get.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ 28+98i }{ 16-196i }\]

OpenStudy (anonymous):

I tried, no idea if its right...

OpenStudy (nubeer):

dont u think 196i^2?

OpenStudy (nubeer):

rather then 196i

OpenStudy (anonymous):

why squared?

OpenStudy (nubeer):

(4+14i)(4-14i) 16 -56i +56i +(14i)(-14i)

OpenStudy (anonymous):

okay...

OpenStudy (nubeer):

do u get now?

OpenStudy (anonymous):

not really

OpenStudy (nubeer):

which part u dont get?

OpenStudy (anonymous):

well what do I do next, I understand the multiplying part what happens next?

OpenStudy (nubeer):

i^2 = -1 right? so -196i^2 = 196.. u get it?

OpenStudy (anonymous):

yes

OpenStudy (nubeer):

so can u solve further nw.?

OpenStudy (anonymous):

I have no clue...Im so sorry...This isnt making any sense

OpenStudy (nubeer):

lol.. which part u have to tell.. \[\frac{ 28+98i }{ 16+196 }\] solve the denominator now .

OpenStudy (anonymous):

212

OpenStudy (nubeer):

\[\frac{ 28 }{ 212 }+\frac{ 98i }{ 212 }\] simplify it

OpenStudy (anonymous):

\[\frac{ 7 }{ 53 } + \frac{ 49 }{ 106 } i\]

OpenStudy (nubeer):

ok thats cool then.. its now ur standard form ..

OpenStudy (anonymous):

lol thanks, sorry it took so long...

OpenStudy (nubeer):

its fine i dont mind.. :)